OpenStudy (brydenwright):

Find the angle between the given vectors to the nearest tenth of a degree. u = <6, -1>, v = <7, -4>

OpenStudy (amistre64):

have you learned about dot product of vectors?

OpenStudy (amistre64):

\[|\vec u|~|\vec v|~cos(\alpha)=\vec u \cdot \vec v \] and solve for \(\alpha \)

OpenStudy (brydenwright):

I tried that, but I got 133.89 and the answer choices are a. 20.3° b. 10.2° c. 0.2° d. 30.3°

OpenStudy (amistre64):

lets see how well you calculated the parts ... what did you get for vector lengths?

OpenStudy (brydenwright):

\[\sqrt{37} and \sqrt{65}\]

OpenStudy (amistre64):

good so sqrt 2405 gives us ...49.0408

OpenStudy (amistre64):

and the dot of the vectors?

OpenStudy (brydenwright):

-1(6)+7(-4)=-34

OpenStudy (amistre64):

<6, -1> <7, -4> 42+4 = 46 is what i get for the dot

OpenStudy (amistre64):

you multiply like parts, and add the results ... your process is a little off is all

OpenStudy (brydenwright):

Oh wow, I made a stupid mistake. Thanks, I can solve it from here!

OpenStudy (amistre64):

good luck :)